eight times a number plus five time another number is -13. The sum of the two numbers is 1. What are the numbers
Answer by macston(5194)

F=first number; S=second number
F+S=1 Subtract S from each side.
F=1-S
8F+5S=-13 Substitute for F.
8(1-S)+5S=-13
8-8S+5S=-13 Subtract 8 from each side
-3S=-21 Divide each side by -3
S=7 ANSWER 1: The second number is 7
F=1-S=1-7=-6 ANSWER 2: The first number is -6.
CHECK
8F+5S=-13
8(-6)+5(7)=-13
-48+35=-13
-13=-13